OpenStudy (theeric):

Hi! I am solving a problem, and I think I correctly arrived at \(\ln\left|e^y-2x\ y\right|=C\). At any rate, I think I'll be left with \(\ln\left|e^y-2x\ y\right|\), and I don't know how to solve for \(y\)!

OpenStudy (theeric):

What I have done is this:

OpenStudy (theeric):

Put each side as the exponent of \(e\). Then \(e^{\ln\left|e^y-2x\ y\right|}=e^C\\\implies e^y-2x\ y=C\) where \(C\) is still just an arbitrary constant.

OpenStudy (anonymous):

Let me guess, you solving a ODE? Not always it is posible to explicitly express y as function of x. Sometimes you just leave a general solution in implicit form

OpenStudy (theeric):

This is from an ODE.... I am supposed to solve it implicitly. Have I done that, then?

OpenStudy (theeric):

Assuming I got there correctly, I mean. I don't want to ask you to check my work.

OpenStudy (anonymous):

yes, you are done

OpenStudy (theeric):

Thank you! :)

OpenStudy (anonymous):

yw